Spectrum:

Hewlp Math:

RodWav3:

wheres question

RodWav3:

oh ok

NoodlesAndRiceYT:

yeah spec i know i said i could help you but nope dis dude can

NineTailedFox:

let me in

NineTailedFox:

please give me acess

Florisalreadytaken:

All you have to do is compare one of the sides -- in \( \triangle PQR\) the opposite is \( 8 \) units, whereas in \( \triangle WXY \) the opposite is \(4\) digits to find the relation between the two of these, (the scale factor) just substitute the values in the following eq. \( \frac{ opposite \ \ of \ \ \triangle WXY }{opposite \ \ of \ \ \triangle PQR} = \frac{8}{4} = \boxed{\frac{2}{1}}\)

Spectrum:

So, it's 2

Spectrum:

I have more,

NineTailedFox:

let me in

Spectrum:

It isn't locked.